A parallel plate capacitor has metal plates, each of area 1.50 m2, separated by 1.40 mm. What charge (in µC) is stored in this capacitor if a voltage of 2.70 103 V is applied to it?
_______µC
Answer #1

first we need to calculate the capacitance of the parallel plate capacitor,

C = (ε * A)/d

where C is the capacitance, A is the area of plate and d is the separation between plates.

C = 8.85 * 10^-12 * 1.5/(1.4 * 10^-3)

C = 9.48 * 10^-9 Farad

Now, expression for the charge stored in a capacitor is,

Q = C*V

Q = 9.48 * 10^-9 * 2.7 * 10^3 C

Q = 25.60 μC

The charge stored in the capacitor is 25.60 μC.